For the reaction

N2(g)+O2(g) = 2NO(g); relate the rate of disappearance of N2 to the rate of formation of NO
Answer #1

Answer

Given chemical reaction is,

N2(g)+O2(g)=2NO(g)

Rate of disappearance of N2 is given as,

Rate of disappearance of N2 = -d[N2]/dt

Rate of formation of NO is given as,

Rate formation of NO = 1/2 d[NO]/dt

Since for the given reaction rate of disappearance of N2 is equal to the twice of rate of formation of NO is given as

Rate= -d[N2]/dt= 1/2 d[NO]/dt